Bill is 6 years younger than his brother Ted. Ed is two years younger than Bill. The sum of the boys' ages is 40. Find Ed's age.
I have tried Ed=x, Ted=y, Bill=x-6 than setting up a system of equation? 2x+y-6=40? I am so confused

Answer by mananth(16946)

Bill is 6 years younger than his brother Ted. Ed is two years younger than Bill. The sum of the boys' ages is 40. Find Ed's age.
Let ted's age be x
bill's age x-6
ed's age = (x-6)-2
..
their sum = 40
x+x-6+x-6-2=40
3x-14=40
add 14 to both sides
3x-14+14=40+14
3x=54
divide by 3
3x/3 = 54/3
x=18 years - Ted's age
..
Bill's age = x-6 18-6= 12 years
..
Ed's age = (x-6)-2 = (18-6)-2= 10 years
CHECK
Their sum = 18+12+10 = 40 years